# Towards Interactive Autonomous Vehicle Testing: Vehicle-Under-Test-Centered Traffic Simulation thanks: This work was supported in part by National Natural Science Foundation of China 52125208, 52232015

### v1

@@ VI Conclusion
The on-road driving safety of autonomous vehicles requires testing under scenarios that closely mirror real-world traffic conditions. In this paper, we propose a Vehicle Under Test (VUT)-Centered environmental Dynamics Inference (VCDI) framework to simulate background traffic objects. The VCDI’s unique capability to integrate a distributional cost function learned from real-world driving data enables the modeling of diverse, strategic responses from surrounding traffic participants to VUT’s maneuvers.
Our findings demonstrate that the VUT-centered augmented input enables background traffic participants driven by DL-based models to dynamically interact with VUT. In addition, the cost-based strategy presets on VUT support directional background traffic simulation, meriting the application of AV testing.

### v2

@@ VI Conclusion
The safety of autonomous vehicles on the road requires rigorous testing in environments that closely mimic real-world traffic conditions. This study introduces a VUT-Centered environmental Dynamics Inference (VCDI) framework to simulate background traffic objects. By leveraging a distributional cost function learned from real-world driving data, the VCDI framework uniquely models diverse, strategic responses from surrounding traffic participants to VUT’s maneuvers.
Our findings reveal that the VUT-centered augmented input enables background traffic participants, driven by DL-based models, to dynamically interact with VUT. In addition, the cost-based strategy presets on VUT facilitate directional background traffic simulation, meriting the application of AV testing.